# HyperMono: A Monotonicity-aware Approach to Hyper-Relational Knowledge Representation

### v2

@@ 4. Experiments
  • RQ1 (Effectiveness): How our model performs compared to the state-of-the-art under different conditions?
@@
  • RQ3 (Parameter sensitivity): How hyper-parameters influence its performance? We present part of the experimental results in the main body, further detailed analysis can be found in Appendix D.
  • RQ4 (Complexity analysis): What is the amount of computation and parameters used by HyperMono?
  • RQ5 (Model transferability): Can HyperMono be used to complete hypergraphs? This results can be found in Appendix E.
@@
4.1. Experiment Setup
4.1.1. Datasets
We conduct experiments on three hyper-relational KGs: WD50K (Galkin et al., 2020), WikiPeople (Guan et al., 2019), and JF17K (Wen et al., 2016), where WD50K and WikiPeople have been extracted from Wikidata (Vrandecic and Krötzsch, 2014) and JF17K from Freebase (Bollacker et al., 2008). We also construct three different sub-datasets, namely: Fixed-percentage Fixed-qualifier, Fixed-percentage Mixed-qualifier, and Mixed-percentage Mixed-qualifier. Statistics and details on the construction can be found in Appendix A.
4.1.2. Baselines.
We compare HyperMono with twelve state-of-the-art baselines for HKGC, including Embedding-based Methods: m-TransH (Wen et al., 2016), RAE (Zhang et al., 2018), NaLP-Fix (Rosso et al., 2020), HINGE (Rosso et al., 2020), and ShrinkE (Xiong et al., 2023); Transformer-based with GNNs Methods: StarE (Galkin et al., 2020) and QUAD (Shomer et al., 2022); Transformer-based without GNNs Methods: Hy-Transformer (Yu and Yang, 2021), GRAN (Wang et al., 2021), HyNT (Chung et al., 2023), HAHE (Luo et al., 2023), and HyperFormer (Hu et al., 2023). Detailed baseline descriptions, evaluation metrics and implementation details can respectively be found in Appendix B and C.
@@ 4.1.1. Datasets
4.2. Main Results
To address RQ1, we conduct experiments on the three types of datasets described above. The corresponding results are shown in Tables 3, 2, and 1.
Fixed-percentage Mixed-qualifier. Table 2 presents the results on WD50K, WikiPeople, and JF17K under different fixed ratios of relational facts with qualifiers. For each dataset, we construct three subsets (as describe in § 4.1.1, Point 2) containing approximately 33%, 66%, and 100% of facts with qualifiers. We observe that under different proportions of hyper-relational knowledge HyperMono also achieves optimal performance. Particularly, for the WikiPeople dataset, in which HyperMono performs poorly in the mixed-percentage mixed-qualifier scenario, adding a certain percentage of qualifiers information to triples results in HyperMono outperforming all baselines. For example, for the Hits@1 metric, HyperMono respectively achieves improvements over the best performing baseline, HyperFormer, of 1.2% / 0.9% / 0.6% in the 33% / 66% / 100% variants. This confirms that the proportion of hyper-relational knowledge is a major factor influencing the performance over WikiPeople under the mixed-percentage mixed-qualifier condition. In contrast, on the WD50K dataset, no significant additional improvement was obtained. The most likely reason is the insufficient richness of hyper-relational knowledge in WD50K. For example, in WD50k_100 65.23% of the triples only contain one hyper-relation pair, which limits the functionality of qualifier monotonicity module.
Table 3. Evaluation of different models under the mixed-percentage mixed-qualifier scenario on the WD50K, WikiPeople and JF17K datasets. All baseline results are collected from the original papers. Best scores are highlighted in bold, the second best scores are highlighted in normal, and ’–’ indicates the results are not reported in previous work. The number in parentheses represents the proportion of triples with hyper-relational knowledge.
Methods WD50K (13.6) WikiPeople (2.6) JF17K (45.9)
MRR Hits@1 Hits@10 MRR Hits@1 Hits@10 MRR Hits@1 Hits@10
m-TransH (Wen et al., 2016) 0.063 0.063 0.300 0.206 0.206 0.463
RAE (Zhang et al., 2018) 0.059 0.059 0.306 0.215 0.215 0.469
NaLP-Fix (Rosso et al., 2020) 0.177 0.131 0.264 0.420 0.343 0.556 0.245 0.185 0.358
HINGE (Rosso et al., 2020) 0.243 0.176 0.377 0.476 0.415 0.585 0.449 0.361 0.624
StarE (Galkin et al., 2020) 0.349 0.271 0.496 0.491 0.398 0.648 0.574 0.496 0.725
GRAN (Wang et al., 2021) 0.489 0.413 0.617 0.591 0.510 0.753
HAHE (Luo et al., 2023) 0.368 0.291 0.516 0.509 0.447 0.639 0.623 0.554 0.806
HyperFormer (Hu et al., 2023) 0.366 0.288 0.514 0.491 0.386 0.654 0.664 0.601 0.787
HyperMono 0.375 0.298 0.522 0.494 0.390 0.657 0.685 0.630 0.796
Mixed-percentage Mixed-qualifier. Table 3 summarizes the results on WD50K, WikiPeople, and JF17K under the mixed-percentage mixed-qualifier condition. We can observe that HyperMono outperforms existing SoTA baselines by a large margin across most metrics on WD50K and JF17K. In particular, when compared with HyperFormer (the best performing baseline), HyperMono respectively achieves 2.1% and 2.9% improvements in MRR and Hits@1 on the JF17K dataset. A similar phenomenon occurred on the WD50K dataset. However, for the WikiPeople dataset, even though we can achieve optimal performance on the Hits@10 metric, we cannot surpass the best performing baseline on MRR and Hits@1. The intuitive reason for this is that, in comparison to the WD50K and JF17K datasets where 13.6% and 45.9% of main triple respectively contain hyper-relational knowledge, only 2.6% of triples in WikiPeople have qualifier pairs. This hinders HyperMono’s ability to exploit stage reasoning and qualifier monotonicity. Specifically, we employ a scaling transformation in the cone space to capture qualifier monotonicity, but the lack of qualifier pairs constrains this functionality.
@@ 4.3. Ablation Studies
4.4. Parameter Sensitivity
To address RQ3, we present some of the parameter sensitivity experiments on WikiPeople-3, including: a) effect of label smoothing; b) effect of embedding dimension $d$. A more detailed analysis of other parameters (i.e., effect of temperature $\gamma$ and effect of initialization range) and other datasets (i.e., WikiPeople-4, JF17K-3, and JE17K-4 datasets) can be found in Appendix D. From Figure 4(a), we observe that the values of label smoothing have an impact on the performance, but there is no evident pattern. However, as shown in Figure 4(b), there is not a strict proportional relationship between embedding dimensions and the actual performance. After reaching a certain value of embedding dimensions, the accuracy stabilizes.
(a) Label Smoothing
(b) Embedding Dimension $d$
Figure 4. Parameter sensitivity experiments under different conditions on Wikipeople-3 dataset.
4.5. Complexity Analysis
To address RQ4, we analyze HyperMono’s complexity on the WD50K, WikiPeople, and JF17K datasets from two perspectives: time complexity and space complexity. The corresponding results are shown in Table 5. The time complexity is measured by the number of floating-point operations (#FLOPs) required during training phase, while the space complexity refers to the amount of a model’s parameters (#Params). The larger the values of #FLOPs and #Params are, the more computing power and higher memory usage are required during the training process. We observe that compared with HyperFormer, HyperMono requires less time and space overhead in each dataset. Specifically, on the WD50K dataset, HyperMono only requires nearly half the #FLOPs and one-sixth the #Params of HyperFormer, which demonstrates that HyperMono is more efficient and can achieve better results while occupying less time and space.
Table 5. Amount of parameters and calculation of the HyperFormer and HyperMono models.
Metrics Model WD50K WikiPeople JF17K
#FLOPs HyperFormer 118.397G 118.167G 118.070G
HyperMono 69.600G 69.371G 69.273G
#Params HyperFormer 66.956M 66.956M 66.956M
HyperMono 10.645M 10.645M 10.645M
5. Conclusion and Future Work
In this paper, we propose HyperMono, a framework for hyper-relational graph completion that properly implements two-stage reasoning and faithfully captures qualifier monotonicity. Specifically, we model each qualifier as a shrinking of the main triple cone to a qualifier cone to realize qualifier monotonicity. Empirical experiments under various settings on the WD50K, WikiPeople, and JF17K datasets demonstrate HyperMono’s robust performance, when compared to competitive baselines. Additionally, we present various ablation studies on parameter sensitivity, complexity analysis and model transferability. For future work, we believe that it is necessary to pay special attention to numerical attribute knowledge in hyper-relational knowledge. How to elegantly integrate schema knowledge into the representation process is also an interesting line of research.
A   Datasets
  1. (1)
  2. (2)
  3. (3)
@@ B   Baselines
We compare HyperMono with twelve state-of-the-art baselines for HKGC, including Embedding-based Methods, Transformer-based with GNNs Methods, and Transformer-based without GNNs Methods:
@@ C   Evaluation Protocol and Implements Details
Implementation Details. We conduct all experiments on six 32G Tesla V100 GPUs, and employ the AdamW (Loshchilov and Hutter, 2019) optimizer and a cosine decay scheduler with a linear warm-up for optimization. The hyperparameters are tuned according to the MRR on the validation set. We use grid search to select the optimal hyperparameters, mainly including: the learning rate $lr\in$3e-4, 4e-4, 5e-4, 6e-4, 7e-4}, the label smooth values $ls\in$0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9}, the embedding dimension $d\in$50, 100, 150, 200, 400, 600}, the temperature of Equation (1) $\gamma\in$1, 2, 3, 4, 5, 6}, the initialization range of embeddings $ir\in$0.01, 0.02, 0.05, 0.1, 0.2, 0.35, 0.5}, the neighbors of the head entity {2, 3, 4}, the number of qualifiers for a relation {5, 6, 7}, the number of layers in the transformer {2, 4, 8}, the number of heads in transformer {1, 2, 4, 8}, the input dropout rate in the transformer {0.6, 0.7, 0.8}.

### v3

@@ 4. Experiments
  • RQ1 (Effectiveness): How our model performs compared to the state-of-the-art under different conditions? We also present the Case study in Appendix H.
@@
  • RQ3 (Parameter sensitivity): How hyper-parameters influence its performance? Additional results can be found in Appendix D and Appendix F.
  • RQ4 (Complexity analysis): What is the amount of computation and parameters used by HyperMono? Results can be found in Appendix E.
  • RQ5 (Model transferability): Can HyperMono be used to complete hypergraphs? Results can be found in Appendix G.
4.1. Experiment Setup
4.1.1. Datasets
We conduct experiments on three hyper-relational KGs: WD50K (Galkin et al., 2020), WikiPeople (Guan et al., 2019), and JF17K (Wen et al., 2016), where WD50K and WikiPeople have been extracted from Wikidata (Vrandecic and Krötzsch, 2014) and JF17K from Freebase (Bollacker et al., 2008). We also construct three different sub-datasets, namely: Mixed-percentage Mixed-qualifier, Fixed-percentage Fixed-qualifier, and Fixed-percentage Mixed-qualifier. Statistics and details on the construction can be found in Appendix A.
4.1.2. Baselines.
We compare HyperMono with fourteen state-of-the-art baselines for HKGC, including Embedding-based Methods: m-TransH (Wen et al., 2016), RAE (Zhang et al., 2018), NaLP-Fix (Rosso et al., 2020), NeuInfer (Guan et al., 2020), HINGE (Rosso et al., 2020), and ShrinkE (Xiong et al., 2023); Transformer-based with GNNs based Methods: StarE (Galkin et al., 2020) and QUAD (Shomer et al., 2022); Transformer-based without GNNs based Methods: Hy-Transformer (Yu and Yang, 2021), GRAN (Wang et al., 2021), HyNT (Chung et al., 2023), HAHE (Luo et al., 2023), HyperFormer (Hu et al., 2023), and HyperCL (Lu et al., 2024). We note that although HELIOS (Lu et al., 2023a), sHINGE (Lu et al., 2023b) and NYLON (Yu et al., 2024) also tackle the HKGC task, there are some variations in terms of experimental design and the selection of evaluation metrics, e.g., HELIOS (Lu et al., 2023a) selects MAP and NDCG as evaluation metrics, sHINGE (Lu et al., 2023b) heavily relies on additional schema knowledge and NYLON (Yu et al., 2024) employs data augmentation techniques. Thus, we do not consider them as baselines. Detailed baseline descriptions, evaluation metrics and implementation details can respectively be found in Appendix B and Appendix C.
@@ 4.1.1. Datasets
4.2. Main Results
To address RQ1, we conduct experiments on the three types of datasets described above. The corresponding results are shown in Table 1, Table 2 and Table 3. As shown in Table 1, the improvement in predicting head and tail entities by HyperMono is less pronounced compared to the improvement on all entities. Consequently, Tables 2 and  3 focus more on the prediction of head and tail entities, which pose greater challenges for HyperMono.
Mixed-percentage Mixed-qualifier. Table 1 summarizes the results on WD50K, WikiPeople, and JF17K under the mixed-percentage mixed-qualifier conditions. We observe that HyperMono outperforms existing SoTA baselines by a large margin across most metrics on WD50K and JF17K in the head/tail entities and all entities scenarios. In particular, in the prediction of head and tail entities, HyperMono respectively achieves 2.1% and 2.9% improvements over HyperFormer (the best baseline) in MRR and Hits@1 on the JF17K dataset. A similar phenomenon occurred on the WD50K dataset. However, for the WikiPeople dataset, even though we can achieve optimal performance on the Hits@10 metric, we cannot surpass the best performing baseline on MRR and Hits@1 in the head/tail and all entities settings. The intuitive reason for this is that, in comparison to the WD50K and JF17K datasets where 13.6% and 45.9% of the main triples respectively contain hyper-relational knowledge, only 2.6% of triples in WikiPeople have qualifier pairs. This hinders HyperMono’s ability to exploit stage reasoning and qualifier monotonicity. Specifically, we employ a scaling transformation in the cone space to capture qualifier monotonicity, but the lack of qualifier pairs constrains this functionality. Furthermore, the proportion of triples with hyper-relational knowledge in JF17K is higher than that in WD50K. We can observe that the performance improvement in JF17K is more significant, which further confirms the importance of the richness of hyper-relational knowledge.
@@ 4.2. Main Results
Fixed-percentage Mixed-qualifier. Table 3 presents the results on WD50K, WikiPeople, and JF17K under different fixed ratios of relational facts with qualifiers. For each dataset, we construct three subsets (cf.  § 4.1.1, Point 2) containing approximately 33%, 66%, and 100% of facts with qualifiers. We observe that under different proportions of hyper-relational knowledge HyperMono also achieves optimal performance. Particularly, for the WikiPeople dataset, in which HyperMono performs poorly in the mixed-percentage mixed-qualifier scenario, adding a certain percentage of qualifiers information to triples results in HyperMono outperforming all baselines. For example, for the Hits@1 metric, HyperMono respectively achieves improvements over the best performing baseline, HyperFormer, of 1.2% / 0.9% / 0.6% in the 33% / 66% / 100% variants. This confirms that the proportion of hyper-relational knowledge is a major factor influencing the performance over WikiPeople under the mixed-percentage mixed-qualifier condition. In contrast, on the WD50K dataset, no significant additional improvement was obtained. The most likely reason is the insufficient richness of hyper-relational knowledge in WD50K. For example, in WD50K_100 65.23% of the triples only contain one hyper-relation pair, which limits the functionality of qualifier monotonicity module.
@@ 4.3. Ablation Studies
5. Conclusions
In this paper, we propose HyperMono, a framework for hyper-relational graph completion that properly implements two-stage reasoning and faithfully captures qualifier monotonicity. Specifically, we model each qualifier as a shrinking of the main triple cone to a qualifier cone to realize qualifier monotonicity. Empirical experiments under various settings and datasets demonstrate HyperMono’s robust performance, when compared to competitive baselines. Additionally, we present various ablation studies on parameter sensitivity, complexity analysis and model transferability. For future work, we believe that it is necessary to pay special attention to numerical attribute knowledge in hyper-relational knowledge. How to elegantly integrate schema knowledge into the representation process is also an interesting line of research.
Appendix
@@ 5. Conclusions
A   Datasets
  1. (1)
  2. (2)
  3. (3)
It is noteworthy that the construction process of the dataset inherently satisfies the qualifier monotonicity property, i.e., for a given main triple, when new qualifier pairs are added to it, the corresponding answer space will not expand. Specifically, the qualifier monotonicity includes two aspects: from nothing to something and from little to more. Taking the WD50K dataset as an example (the same applies to any other dataset), the content prefix with P and Q below corresponds to the ID value from the WikiData knowledge graph.
  • From nothing to something: This phrase encapsulates the process of extending a main triple into a hyper-relational fact by incorporating qualifier pairs. For instance, given the main triple (Q200405, P1411, ?), the possible answers to ? are Q4834546, Q593098, and Q103916. By introducing a qualifier pair $\mathcal{Q}_{1}$ = (P1868: Q272860), the resulting hyper-relational fact is defined as $q_{1}$ = {(Q200405, P1411, ?), (P1868: Q272860)}. Consequently, the set of valid answers to ? is refined to Q593098 and Q103916, effectively eliminating Q4834546 as a possibility. This demonstrates how the incorporation of qualifiers progressively constrains the answer space. During dataset construction, we begin by extracting the main triple and iteratively introduce relevant qualifier pairs, gradually refining the possible answer set. The concept of from nothing to something signifies a fundamental transition from conventional triple-based representations to hyper-relational facts enriched with hyper-relational knowledge, marking a significant conceptual shift from 0 to 1.
  • From little to more refers to the progressive enrichment of hyper-relational facts through the incorporation of additional qualifier pairs. For instance, by appending the qualifier pair $\mathcal{Q}_{2}$=(P805: Q767355) to $q_{1}$, we obtain $q_{2}$={(Q200405, P1411, ?), (P1868: Q272860), (P805: Q767355)}. Consequently, the corresponding answer to ? is refined to Q103916, thereby further narrowing the answer set by excluding Q593098. This process exemplifies the transition from an incomplete hyper-relational knowledge representation to a more comprehensive one, enhancing the specificity and informativeness of the captured knowledge.
@@ B   Baselines
We compare HyperMono with fourteen state-of-the-art baselines for HKGC, including Embedding-based Methods, Transformer-based with GNNs Methods, and Transformer-based without GNNs Methods:
@@
  • NeuInfer (Guan et al., 2020) distinguishes the information in the same n-ary fact and represented each n-ary fact as a primary triple coupled with a set of its auxiliary descriptions.
@@
  • HyperCL (Lu et al., 2024) designs relation-aware graph attention networks to capture the hierarchical ontology and a concept-aware contrastive loss to alleviate the dominance issue.
@@ C   Evaluation Protocol and Implements Details
Implementation Details. We conduct all experiments on six 32G Tesla V100 GPUs, use the Pytorch 1.8.1 framework (it should be mentioned that using Pytorch 2.1.2 will result in incorrect running results), and employ the AdamW (Loshchilov and Hutter, 2019) optimizer and a cosine decay scheduler with a linear warm-up for optimization. The hyperparameters are tuned according to the MRR on the validation set. We use grid search to select the optimal hyperparameters, mainly including: the learning rate $lr\in$3e-4, 4e-4, 5e-4, 6e-4, 7e-4}, the label smooth values $ls\in$0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9}, the embedding dimension $d\in$50, 100, 150, 200, 400, 600}, the temperature of Equation (1) $\gamma\in$1, 2, 3, 4, 5, 6}, the initialization range of embeddings $ir\in$0.01, 0.02, 0.05, 0.1, 0.2, 0.35, 0.5}, the neighbors of the head entity {2, 3, 4}, the number of qualifiers for a relation {5, 6, 7}, the number of layers in the transformer {2, 4, 8}, the number of heads in transformer {1, 2, 4, 8}, the input dropout rate in the transformer {0.6, 0.7, 0.8}. For the reproducibility of the model, we also provide the running logs of HyperMono in different scenarios in the anonymous code link: https://anonymous.4open.science/r/HKGC-HyperMono-AE61.